OpenStudy (anonymous):

Solve -2(2x + 5) - 3 = -3(x - 1) 16 -16 -10 -13 I'm not sure if it's 16 or -16. I solved it twice and got both of those answers. Please help Immediately

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) I saw you ask for help through the chat... distribute so -4x - 10 - 3 = -3x + 3 add like terms -4x - 13 = -3x + 3 add 4x to both sides -13 = -3x + 4x + 3 subtract 3 from both sides -13 - 3 = x x = -13 - 3 = ?

OpenStudy (anonymous):

Instead of adding 4x to both sides, I added 3x to both sides. Then I got -x - 13 = 3. The I added 13 to each side and got -x = 16. So my final answer would be x = -16, right?

OpenStudy (sleepyhead314):

correct! :)

OpenStudy (anonymous):

Thank you!

OpenStudy (sleepyhead314):

Glad I was able to help ^_^